OpenStudy (kamille):

Hello, I need to find \[A ^{2}-B ^{2}\] if \[A=\frac{ 1 }{ 2}(2^{x}+2^{-x)}\] \[B=\frac{ 1 }{ 2 }(2^{x}-2^{-x})\] Any tips?

OpenStudy (kamille):

@Mertsj

hartnn (hartnn):

\(A^2-B^2= (A+B)(A-B)\) first, find A+B= ... ?

OpenStudy (kamille):

do I need, firstly, "unbracket" it? Like \[2^{-1}(2^{x}+2^{-x})=2^{x-1}+2^{-x-1}\] or not?

hartnn (hartnn):

no need, \((1/2) [2^x+2^{-x}] +(1/2) [2^x-2^{-x}] = (1/2) [2^x+2^{-x} +2^x-2^{-x} ]=...? \)

OpenStudy (kamille):

is it? \[ 2*2^{x}=2^{x+1}\]

OpenStudy (kamille):

i have forgotten 1/2

OpenStudy (kamille):

then it is 2^x

OpenStudy (kamille):

rigth or not?

hartnn (hartnn):

yes, A+B = 2^x is correct. what about A-B , similarly ?

OpenStudy (kamille):

2^-x?

hartnn (hartnn):

good! absolutely correct :)

hartnn (hartnn):

so, what about (A+B)(A-B) = 2^x * 2^(-x) =... ?

OpenStudy (kamille):

2^x+(-x)=2^x-x=2^0=1?

OpenStudy (kamille):

tank you a lot, hartnn! You have helped me a lot!!

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\) welcome ^_^